OpenStudy (anonymous):

You start from your home, move 5 miles north, then 10 miles east, then 8 miles south, and finally 14 miles west. How long is the straight line from your final location to your home?

OpenStudy (anonymous):

@iambatman im sure he knows that he can drow a picture but thats not help lol

OpenStudy (anonymous):

yeah i concur

OpenStudy (anonymous):

@iambatman uhhhh

OpenStudy (anonymous):

soooo..can i get some help, or nah?

OpenStudy (anonymous):

@Hood_Problmz nah! lol i wish i knew but i can try to help you figuire it out or nah? lmaoo

OpenStudy (anonymous):

@Hood_Problmz yay thats ur answer ! (:

OpenStudy (anonymous):

damn. aite sure whatever you bring to the table is wonderful

OpenStudy (anonymous):

aite batman my grades are depending on this one answer let me type this into the test and see if it's correct...

OpenStudy (anonymous):

@Hood_Problmz but he might not be correct or he might your choice to take his answer

OpenStudy (anonymous):

Dammit @iambatman dat pellet was not correct, but aye thnx for trying family.

OpenStudy (anonymous):

awwwww

OpenStudy (anonymous):

@Hood_Problmz what was ur grade?

OpenStudy (anonymous):

it's just #s not directions

OpenStudy (anonymous):

LMMFAO

OpenStudy (anonymous):

@iambatman you stupid ! frfr

OpenStudy (anonymous):

i fail if i get less than 4 outta 5 correct

OpenStudy (anonymous):

@iambatman no im serious u r

OpenStudy (anonymous):

@Hood_Problmz grade?

OpenStudy (anonymous):

f

OpenStudy (anonymous):

@iambatman so u tricked him?? lol

OpenStudy (anonymous):

muthaflutteraa...

OpenStudy (anonymous):

@iambatman GTFO !

OpenStudy (anonymous):

one sec yall imma restart da test stay online batman don't come wit dat bull ish mane!!!!!!!!

OpenStudy (anonymous):

What is the equation of a circle with its center at (-h, k) and radius r.?

OpenStudy (anonymous):

anybody there?

OpenStudy (anonymous):

@keke_luvs_u where u guh?!

OpenStudy (anonymous):

@Hood_Problmz im here lmao

OpenStudy (anonymous):

(x-h)^2+(y-k)^2 = r^2 now it's up to you to believe me if I'm right or wrong :)

OpenStudy (anonymous):

@iambatman don't get shy on me now maneee!!!

OpenStudy (anonymous):

das fuke?

OpenStudy (anonymous):

\[(x-h)^2+(y-k)^2=r^2\]

OpenStudy (anonymous):

oooooooo killem!!!!!!!!!!!

OpenStudy (anonymous):

i got dat ish right!

OpenStudy (anonymous):

:P

OpenStudy (anonymous):

yay yay ! what was grade??

OpenStudy (anonymous):

What are the center and radius of the given circle? x^2 – 12x + y^2 – 16y = 0

OpenStudy (anonymous):

aw it ain't ova yet dats just 1

OpenStudy (anonymous):

@iambatman mane fk dis sh*t cuh!

OpenStudy (anonymous):

@Hood_Problmz ??

OpenStudy (anonymous):

oh god, here we go

OpenStudy (anonymous):

what?

OpenStudy (anonymous):

he gon lecture me

OpenStudy (anonymous):

@Hood_Problmz what grade r u??

OpenStudy (anonymous):

like my grade in school?

OpenStudy (anonymous):

@Hood_Problmz ya

OpenStudy (anonymous):

my bad my momz rolled up oon me and i ahd to click off

OpenStudy (anonymous):

im a junior in high school

OpenStudy (anonymous):

@Hood_Problmz what grade r u !!?? not school

OpenStudy (anonymous):

I'm out peace

OpenStudy (anonymous):

-______-...f....again

OpenStudy (anonymous):

aite bat preciate cuz

OpenStudy (anonymous):

@Hood_Problmz death or cant read which one??

OpenStudy (anonymous):

@keke_luvs_u well damn...

OpenStudy (anonymous):

neither ig

OpenStudy (anonymous):

@Hood_Problmz then answer !!!!!!

OpenStudy (anonymous):

@keke_luvs_u LOL i said f like 5 years ago

OpenStudy (anonymous):

@Hood_Problmz bye

OpenStudy (anonymous):

@keke_luvs_u damn what did i say?

OpenStudy (anonymous):

@Hood_Problmz u didnt answer my question...what grade are you?

OpenStudy (anonymous):

I got a f

OpenStudy (anonymous):

@Hood_Problmz your a idoit !! i meant what grade are you 8th 9th what !!!!!!!!!

OpenStudy (anonymous):

damn i said 11th and then you said not that type of grade!

OpenStudy (anonymous):

whoops lol ok ok

OpenStudy (anonymous):

@keke_luvs_u lol u silly